OpenStudy (anonymous):

Louisa walked 3 and 1/4 of a mile on Monday, 2 and 4/5 mile on Tuesday, and 1 and 1/8 mile today. ESTIMATE the total distance she walked on these 3 days. Thank you and could you explain how you did it?.

OpenStudy (hunus):

The total distance she walked is going to be the sum of all of the distances. The first day she walked 3 1/4 of a mile, the second day she walked 2 4/5 of a mile and the third day she walked 1 1/8 of a mile. total distance = 3 1/4 + 2 4/5 + 1 1/8 We can add up the whole number portions of these number and save the fraction part for after. total distance = 3 + 2 + 1 + 1/4 + 4/5 + 1/8 = 6 + 1/4 + 4/5 + 1/8 Now we have to add the fraction portions. which is 1/4 + 4/5 + 1/8 to do this we need to find a common denominator for all of these numbers (unless we were to convert them to decimals) The tricky number here is going to be finding the multiple of 8 that goes with both 4 and 5. I'm going to try 8 * 5 (because I know both 8 and 5 goes in to it) and I get 40 which 4 goes into also. 1/4 + 4/5 + 1/8 = 10/40 + 32/40 + 5/40 = 47/40 = 1 7/40 Now we can add this to our whole part total distance = 6 + 1 7/40 = 7 7/40

OpenStudy (anonymous):

just add 3 1/4 , 2 4/5 and 1 1/8. You can convert them into improper fractions - 13/4 + 14/5 + 9/8 The L.C.M of 4,5,8 = 40. so we have - 13x10/4x10 + 14x8/5x8 + 9x5/8x5 (10,8,5 is the factor which makes 4,5,8 respectively 40 on multiplying) +> 130+112+45/40 => 287/40 = approx 7.175 miles

OpenStudy (anonymous):

thank you Hunus and luckythebest

OpenStudy (hunus):

Yup :)